The question ask: The graph of y=2x^2+12x+1 is: a parabola whose vertex is ?
Answer by Alan3354(69443)

The graph of y=2x^2+12x+1 is: a parabola whose vertex is ?
---------------
It's at x = -b/2a
x = -12/4 = -3
f(-3) = -17
V at (-3,-17)